# Identification of Single-Treatment Effects in Factorial Experiments

### v2

@@ 2.2 Sufficient parametric assumptions for point identifiability
If one is willing to make parametric assumptions, then single-treatment quantities can be identifiable using factorial and observational data. For instance, when the data-generating process satisfies linearity with interactions, $y_{i}=\beta_{0}+\beta_{1}a_{i}+\beta_{2}b_{i}+\beta_{3}a_{i}b_{i}+u_{i}$, it can be shown that (proof in Theorem C.2):
@@
$ATE=\sum_{b\ \in\ \operatorname{supp}(B)}(\operatorname{\mathbb{E}}(Y_{a,b})-\operatorname{\mathbb{E}}(Y_{a^{\prime},b}))\Pr(B=b)$ (2)
To the best of my knowledge, the solution that combines factorial and observational data was initially proposed by De la Cuesta et al., (2022). This estimand aligns with the Average Marginal Component Effect (AMCE) described by Hainmueller et al., (2014), where factorial quantities are weighted by the population distribution of $B$. That case was also studied by Zhao and Ding, (2022).
In some cases, even when single-treatment quantities are not identifiable, it is still possible to identify their Average Treatment Effect (ATE) by making the additive separability (A.S.) assumption. This assumption posits that $y_{i}=f(a_{i},u_{i})+g(b_{i},u_{i})$, where $f$ and $g$ are arbitrary functions. Since this functional form implies $\operatorname{\mathbb{E}}(Y_{a,b})-\operatorname{\mathbb{E}}(Y_{a^{\prime},b})=f(a,u_{i})-f(a^{\prime},u_{i})$ is constant for any $b$, the assumption is also referred to as ”no interaction.” This implies that the ATE is identifiable through:
@@
$\sum_{b}(\operatorname{\mathbb{E}}(Y_{a_{1},b})-\operatorname{\mathbb{E}}(Y_{a_{0},b}))\cdot\sigma(b),\quad\text{where $\sigma(b)$ is the uniform p.m.f. over $B$}.$ (3)
This estimand aligns with the Average Marginal Component Effect (AMCE) described by Hainmueller et al., (2014), where factorial quantities are weighted by a uniform distribution of $B$. The formal proofs supporting this assertion are provided in Theorem C.1 in the appendix.
@@
It is important to emphasize that although the Average Treatment Effect (ATE) may be identifiable, the ”no interaction” assumption does not guarantee the identifiability of single-treatment expected outcomes, such as $\operatorname{\mathbb{E}}(Y_{a})$, even in the presence of both factorial and observational data. This outcome is demonstrated in Theorem B.2. This situation arises because when computing the ATE, the constant component implied by the absence of interaction is subtracted from itself, but that does not happen when computing the simple expected outcomes.
@@
Lastly, assuming a strictly linear model, expressed as $y_{i}=\beta_{0}+\beta_{1}a_{i}+\beta b_{i}+u_{i}$, enables the direct utilization of both interactive linearity and additive separability. As a result, single-treatment quantities as well as the Average Treatment Effect (ATE) become point identifiable. This conclusion is concisely encapsulated in Corollary C.2.1.
@@
In summary, adopting the assumptions of additive separability and linear interactions may facilitate the identification of single-treatment quantities or risk differences. However, these assumptions should not be considered innocuous. For instance, assuming additive separability implies the absence of interaction between treatments, suggesting that the effect of one treatment remains unchanged in the presence of another—a potentially overly restrictive assumption in certain scenarios. In Section 4, I will present a method for conducting sensitivity analysis when restricting the number of interactive units in the experiment. Furthermore, the assumption of linearity may be overly constraining.222 It is important to note that when referring to linearity, we imply the linearity of the real-world model—a crucial identification assumption. This does not necessarily imply the use of linear regression for estimation, as linear unbiased estimators can be utilized even in non-linear settings. Particularly when dealing with binary or discrete outcomes, assuming linearity could result in overly restrictive models, effectively constraining the analysis within vector spaces over fields $F_{0,1}$. Nevertheless, these parametric assumptions are merely illustrative, and researchers are not precluded from justifying other, less stringent assumptions that are sufficient for identification.
2.3 Sufficient structural assumptions for point identifiability
F-bias is typically problematic when one is not willing to make parametric assumptions. There are cases when making structural and/or functional assumptions might be justifiable by researchers. If one is willing to relax confounding of some variables, then the problem is eliminated.
There are three typical cases where single-treatment effects are identifiable given structural assumptions (non-confounding). In Figure 4-a, for instance, $B$ can be confounded with $Y$, but there is no effect from $B$ on $Y$. In cases represented by this DAG, then $\Pr(Y_{a,b}=y)=\Pr(Y_{a}=y)$, as interventions on $B$ do not cause differences in $Y$. So, independently of the distribution of $B$ employed to average $\Pr(Y_{a,b}=y)$, $\Pr(Y_{a,b}=y)=\Pr(Y_{a}=y)$. This example is relevant because it justifies adding new attributes and interventions in experiments just to improve realism. Without this justification, one might fear that even irrelevant interventions might imply biased estimates, but indeed they do not. For instance, a researcher might be concerned that interventions, such as the shape of a pill (rounded or squared), could impact experiments about the effect of that pill on health outcomes. Indeed, because the shape of a medication is not thought to causally influence the outcome, they might be ignored – even if in reality they might be confounded with the outcome.
A second case is represented in Figure 4-b. Here $B$ indeed causes $Y$, but they are unconfounded. In this case, then one can apply rule-2 of do-calculus (Pearl,, 1995) to show that $\Pr(Y_{a,b}=y)=\Pr(Y_{a}=y|B=b)$. For this example, if we have the distribution of $B$ in the population, $\Pr(b)$, then $\Pr(Y_{a,b}=y)=\Pr(Y_{a}=y\ |\ B=b)$. So,
Finally, the third scenario is represented in Figure 4-c and can be understood as a type of the DAG ’b’ case. Here, no variable is confounded with Y. Then the effect $\Pr(Y_{a}=y)$ can be immediately identifiable by Equation 3,i.e. assuming a uniform distribution of $B$. Therefore, only factorial data would be sufficient for identifying the single-treatment quantity. This third scenario corresponds to a case where the goal is to explore the effects of two programs without investigating the causal factors behind them. For instance, Duflo et al., (2015) tested the effect of two educational programs in Kenya on the probability of students dropping out of school, having an early marriage, or becoming pregnant. Those two programs are 1) educational subsidies, including free school uniforms, and 2) a national HIV curriculum to be taught to students. Because those programs are often interventional and cannot be confounded, then the single-treatment effect of each can be directly estimated without bias. However, if one intends to estimate not the effect of each program, but the effect of HIV education or certain subsidies in general, then researchers have to be more careful. The difference is subtle but is relevant for identifiability and estimation.
$B$$Y$$A$(a)$B$$Y$$A$(b)$B$$Y$$A$(c)
@@ 2.3 Sufficient structural assumptions for point identifiability
Table 1: Identification results with respect to parametric assumptions. Structurally, it is assumed that the graph in Figure 3 holds. Generally, identification of single-treatment quantities is not feasible in such settings, even if monotonicity of both treatments is assumed. However, if researchers possess only factorial or observational distributions individually, identification of single-treatment probabilities as well as the Average Treatment Effect (ATE) is possible only if they can justify that the real-world model is linear without interactions. Otherwise, this is achievable only when researchers have both factorial and observational distributions in the interactive linear case. Finally, if the assumption of no interaction holds, the ATE can be identified using both distributions, but single-treatment probabilities cannot.
Data Estimand General (+/- Monotonic) No Interaction (+/- Monotonic) Linearity + Interaction Linearity + No Interaction
@@
Table 2: Identification results of $\operatorname{\mathbb{E}}[Y_{a}]$ with respect to structural assumptions. Columns indicate which distributions are sufficient for identification (only observational, only factorial, or both). When necessary, cells indicate which rule of do-calculus enable identification, or if confounding prevents it. Nonparametrically (Fig. 3), neither factorial nor observational distributions allow for the identification of single-treatment results (no assumption). If one assumes that $B$ does not cause $Y$ (Fig. 4-a), then only the factorial distribution is adequate for identification, as per rule-3 of do-calculus. If one assumes that $B$ and $Y$ are not confounded (Fig. 4-b), then both factorial and observational data are necessary for identification. Finally, if neither $B$ nor $A$ are confounded with $Y$ (Fig. 4-c), then observational and factorial data become equivalent, leading to identification in both cases.
@@
Unconfounded $A$,$B$, and $Y$ Fig. 4-c Yes (rule 1) Yes (rule 3) Yes (rule 3)
@@ 3.1 Closed-form bounds
Up to this point, I have highlighted the challenges in identifying single-treatment quantities from factorial experiments. I propose two solutions: a) partial identification, and b) sensitivity analysis. This section focuses on the partial identification solution, deriving sharp bounds for the desired quantities. The next section will address the sensitivity analysis solution.
@@
In order to calculate the sharp bounds, I use the same procedure as the algorithm autobounds, derived by Duarte et al., (2023). This approach consists in reducing causal questions with discrete data to polynomial programming problems, for which it is possible to sharply bound causal effects using efficient dual relaxation and spatial branch-and-bound techniques. From the perspective of an applied researcher, autobounds is straightforward to use. The user (1) provides a causal diagram describing the process under study; (2) states assumptions; (3) states the quantity of interest; and (4) inputs available data, however incomplete or mismeasured. The algorithm then outputs the most precise possible answer given these parameters—which may be a range of indistinguishable, observationally equivalent possibilities. Thus, the method fully automates the process of computing causal bounds—i.e., rigorously identifying all possible answers that cannot be ruled out by an observed dataset. These results are made possible by the theoretical work proving that any causal problem can be transformed into an equivalent polynomial programming problem. Then, it becomes straightforward to derive a general algorithm for doing the problem reduction, which opens the door to a vast range of highly efficient polynomial programming solution techniques such as spatial branch-and-bound and linear programming relaxations. This method is an extension of the original Balke and Pearl, (1997)’s linear programming technique and employs recent results in causality (Evans,, 2018; Rosset et al.,, 2018).
@@
In the appendix, I have also derived bounds for estimands such as $\operatorname{\mathbb{E}}(Y_{a})$ and the ATE in two other scenarios: one with only factorial data, and another with both observational and factorial data, under the monotonicity assumption. There are clear limitations to those scenarios for cases involving treatments with more than two levels and other causal structures. However, these limitations can be circumvented by employing the provided code, which makes use of autobounds.
Despite being bounded in scenarios with binary outcomes, these bounds can be extended to the discrete case efficiently. One simply needs to compute bounds for each possible outcome of $y$. For instance, if $Y$ has three levels $y_{0},y_{1},y_{2}$, one calculates the bounds three times: first for $\Pr(Y_{a}=y_{0})$, then for $\Pr(Y_{a}=y_{1})$, and $\Pr(Y_{a}=y_{2})$. During each bound computation, the other two levels are aggregated against the main one. When the outcome is continuous, a useful strategy suggested by both Balke and Pearl, (1997) and Levis et al., (2023) is to calculate causal bounds for $\Pr(Y_{a}<k)$, for all $k$ within the bounded range of $Y$. In practice, this approach implies some form of discretization, ensuring valid bounds even if they are no longer sharp.
@@ 3.2 Sensitivity Analysis
The concept behind this sensitivity analysis acknowledges the unrealistic nature of the ”no interaction” assumption. Instead of outright accepting this, one could restrict the proportion of units exhibiting some level of interaction. For example, a parameter like $\theta=0.2$ acknowledges the potential presence of interactive units while also indicating that they may not exceed 20% of the population. Additionally, one can determine the maximum proportion of interactive units that could nullify the ATE or push it towards a specific value. Therefore, this procedure should be applied whenever presenting results on single-treatment effects, particularly with factorial data.

### v3

@@ 2.2 Sufficient parametric assumptions for point identifiability
If one is willing to make parametric assumptions, then single-treatment quantities can be recovered using factorial and observational data. For instance, when the data-generating process satisfies linearity with interactions, $y_{i}=\beta_{0}+\beta_{1}a_{i}+\beta_{2}b_{i}+\beta_{3}a_{i}b_{i}+u_{i}$, it can be shown that (proof in Theorem C.2):
@@
$ATE=\sum_{b\ \in\ \operatorname{supp}(B)}(\operatorname{\mathbb{E}}(Y_{a,b})-\operatorname{\mathbb{E}}(Y_{a^{\prime},b}))\Pr(B=b),$ (2)
where $P(B=b)$ is the population distribution of $B$. This linearity is not mere linearity of Conditional Expectation Function (CEF), but rather linearity of the real-world model, which is a stronger assumption.
This estimand remains extremely relevant in the conjoint experiment literature, as it aligns with the Average Marginal Component Effect (AMCE) described by Hainmueller et al. (2014), where factorial quantities are weighted by the population distribution of $B$. That case was also studied by Zhao and Ding (2022). In other words, the ATE is recoverable by weighting the factorial effects with the population distribution of $B$, which is exactly the AMCE. Notice, however, that $P(B=b)$ is the population distribution of $B$, not merely the uniform distribution of $B$ in the factorial experiment. Although the equation appears just a mere application of the law of total probability, in fact, it requires data fusion methods to estimate it, as $P(B=b)$ is not directly observable in the factorial experiment. This solution was initially proposed in De la Cuesta et al. (2022) under the name Population Average Marginal Component (pAMCE).
On the other hand, if the linearity form assumption is too restrictive, it is still possible to identify their ATE by making the additive separability (AS) assumption. This assumption posits that $y_{i}=f(a_{i},u_{i})+g(b_{i},u_{i})$, where $f$ and $g$ are arbitrary functions. Since this functional form implies $\operatorname{\mathbb{E}}(Y_{a,b})-\operatorname{\mathbb{E}}(Y_{a^{\prime},b})=f(a,u_{i})-f(a^{\prime},u_{i})$ is constant for any $b$, the assumption is also referred to as “no interaction” or “homogeneity.” This implies that the ATE is identifiable through:
@@
$\sum_{b}(\operatorname{\mathbb{E}}(Y_{a_{1},b})-\operatorname{\mathbb{E}}(Y_{a_{0},b}))\cdot h(b),\quad\text{where $h(b)$ is the uniform p.m.f. over $B$}.$ (3)
In other words, under the AS assumption, the single-treatment ATE can be recovered using either any fixed level of $B$ or by averaging over $B$ with a uniform distribution. That suggests that AS or no interaction is sufficient for external validity of the ATE from factorial experiments to single-treatment settings. The formal proofs supporting this assertion are provided in Theorem C.1 in the appendix.
@@
It is important to emphasize that although the ATE may be recovered, the AS assumption does not guarantee the identifiability of single-treatment point exposures, even in the presence of both factorial and observational data. In other words, despite $\operatorname{\mathbb{E}}(Y_{a}-Y_{a^{\prime}})$ being identifiable, $\operatorname{\mathbb{E}}(Y_{a})$ itself remains unidentifiable. This outcome is demonstrated in Theorem B.2. This situation arises because when computing the ATE, the constant component implied by the absence of interaction is subtracted from itself, but that does not happen when computing the simple expected outcomes.
@@
Lastly, if one is willing to make both the AS assumption and linearity (non-interactive linearity), expressed as $y_{i}=\beta_{0}+\beta_{1}a_{i}+\beta b_{i}+u_{i}$, then one obtain the benefits of both assumptions. As a result, single-treatment point exposures as well as the ATE become point identifiable. This conclusion is concisely encapsulated in Corollary C.2.1.
@@
In summary, adopting the assumptions of AS and linear interactions may facilitate the identification of single-treatment point exposures or risk differences. However, these assumptions should not be considered innocuous. For instance, assuming AS implies the absence of interaction between treatments, suggesting that the effect of one treatment remains unchanged in the presence of another—a potentially overly restrictive assumption in certain scenarios. In Section 4, I will present a method for conducting sensitivity analysis when restricting the number of interactive units in the experiment. Furthermore, the assumption of linearity may be overly constraining. Particularly when dealing with binary or discrete outcomes, assuming linearity could result in overly restrictive models, effectively constraining the analysis within vector spaces over fields $F_{0,1}$. Nevertheless, these parametric assumptions are merely illustrative, and researchers are not precluded from justifying other, less stringent structural assumptions that are sufficient for identification. All these results are summarized in Table 1.
@@
Table 1: Point-identification results with respect to parametric assumptions. Structurally, it is assumed that the graph in Figure 3 holds. Generally, identification of single-treatment quantities is not feasible in such settings. However, if researchers possess only factorial or observational distributions individually, identification of single-treatment probabilities as well as the ATE is possible only if they can justify that the real-world model is linear without interactions. Otherwise, this is achievable only when researchers have both factorial and observational distributions in the interactive linear case. Finally, if the assumption of no interaction holds, the ATE can be identified using both distributions, but single-treatment probabilities cannot.
Data Estimand General AS (No Interaction) Linearity + Interaction Linearity + AS
@@
2.3 Sufficient structural assumptions for point identifiability
F-bias is typically problematic when one is not willing to make parametric assumptions. However, there are cases when making structural and/or functional assumptions might be justifiable by researchers. If one is willing to relax confounding of some variables in the single-treatment target population, then the problem is eliminated. In fact, the literature has already developed methods to deal with the problem of identifying causal quantities from multiple data sources with different interventional patterns (Bareinboim and Pearl, 2016; Lee et al., 2020; Lee and Shpitser, 2020). Here I illustrate three typical cases where single-treatment quantities are identifiable given structural assumptions (non-confounding).
There are three typical cases where single-treatment effects are identifiable given structural assumptions (non-confounding) over the target population. In Figure 4-a, for instance, $B$ can be confounded with $Y$, but there is no effect from $B$ on $Y$. In cases represented by this graph, then $\Pr(Y_{a,b}=y)=\Pr(Y_{a}=y)$. [^4] This does not imply that $\Pr(Y_{a}=y|B=b)=\Pr(Y_{a}=y)$. Confusing causal interaction with effect modification is a conceptual blunder. So, independently of the distribution of $B$ employed to average $\Pr(Y_{a,b}=y)$, $\sum_{b\ \in\ \operatorname{supp}(B)}\Pr(Y_{a,b}=y)P(B=b)=\Pr(Y_{a}=y)$. This example is relevant because it justifies adding new attributes and interventions in experiments just to improve realism. Without this justification, one might fear that even irrelevant interventions might imply biased estimates, but indeed they do not. For instance, a researcher might be concerned that interventions such as whether a medication is prescribed as a brand-name or a generic formulation could affect studies of that medication’s impact on health outcomes. Because brand and generic formulations contain the same active ingredient and are designed to be therapeutically equivalent, such differences are often assumed not to causally influence the outcome and are therefore ignored. However, in practice, brand versus generic status may be correlated with patient characteristics, insurance coverage, prescribing settings, or health-care access, all of which are themselves associated with the outcome. As a result, brand status can be confounded with the outcome even in the absence of a direct causal effect.
A second case is represented in Figure 4-b. Here $B$ indeed causes $Y$, but they are unconfounded. In this case, then one can apply rule-2 of do-calculus (Pearl, 1995) to show that $\Pr(Y_{a,b}=y)=\Pr(Y_{a}=y|B=b)$, i.e. causal interaction is equivalent to effect modification. For this example, if we have the distribution of $B$ in the population, $\Pr(b)$, then $\Pr(Y_{a,b}=y)=\Pr(Y_{a}=y\ |\ B=b)$. So,
Finally, the third scenario is represented in Figure 4-c and can be understood as a type of the DAG ‘b’ case. Here, no variable is confounded with Y. Then the effect $\Pr(Y_{a}=y)$ can be identified by Equation 3, i.e. assuming a uniform distribution of $B$. Therefore, factorial data alone would be sufficient for identifying the single-treatment quantity. This third scenario corresponds to a case where the goal is to explore the effects of two programs without investigating the causal factors behind them. For instance, Duflo et al. (2015) tested the effect of two educational programs in Kenya on the probability of students dropping out of school, having an early marriage, or becoming pregnant. Those two programs are 1) educational subsidies, including free school uniforms, and 2) a national HIV curriculum to be taught to students. Because those programs are often interventional and cannot be confounded, then the single-treatment effect of each can be estimated without additional adjustment. However, if one intends to estimate not the effect of each program, but the effect of HIV education or certain subsidies in general, then researchers have to be more careful. The difference is subtle but is relevant for identifiability and estimation. This third situation is relevant in the scientific practice, because although the evaluation of a program might be successful, the ultimate goal of understanding the effect of a particular factor on a general population might not be achieved, especially in presence of confounding with other factors.
BBYYAA(a)BBYYAA(b)BBYYAA(c)
@@ 2.3 Sufficient structural assumptions for point identifiability
Table 2: Point-identification results of $\operatorname{\mathbb{E}}(Y_{a})$ with respect to structural assumptions. Columns indicate which distributions are sufficient for identification (only observational, only factorial, or both). When necessary, cells indicate which rule of do-calculus enable identification, or if confounding prevents it. Nonparametrically (Fig. 3), neither factorial nor observational distributions allow for the identification of single-treatment results (no assumption). If one assumes that $B$ does not cause $Y$ (Fig. 4-a), then only the factorial distribution is adequate for identification, as per rule-3 of do-calculus. If one assumes that $B$ and $Y$ are not confounded (Fig. 4-b), then both factorial and observational data are necessary for identification. Finally, if neither $B$ nor AA are confounded with $Y$ (Fig. 4-c), then observational and factorial data become equivalent, leading to identification in both cases.
@@
Unconfounded AA,$B$, and $Y$ Fig. 4-c Yes (rule 1) Yes (rule 3) Yes (rule 3)
@@ 3.1 Closed-form bounds
Up to this point, I have highlighted the challenges in identifying single-treatment quantities from factorial experiments. In particular, I showed that some assumptions are sufficient for point-identification, but they might be too restrictive in practice. If those assumptions are thought to fail, or if researchers are in doubt about their validity, then point-identification is not achievable.
When these assumptions are not justifiable, I propose two solutions: a) partial identification, and b) sensitivity analysis. This section focuses on the partial identification solution, deriving sharp bounds for the desired quantities. The next section will address the sensitivity analysis solution.
@@
In order to calculate the sharp bounds, I use the same procedure as the algorithm autobounds, derived by Duarte et al. (2023). This approach consists in reducing causal questions with discrete data to polynomial programming problems, for which it is possible to sharply bound causal effects using efficient dual relaxation and spatial branch-and-bound techniques. From the perspective of an applied researcher, autobounds is straightforward to use. The user (1) provides a causal diagram describing the process under study; (2) states assumptions; (3) states the quantity of interest; and (4) inputs available data, however incomplete or mismeasured. The algorithm then outputs the most precise possible answer given these parameters—which may be a range of indistinguishable, observationally equivalent possibilities. Thus, the method fully automates the process of computing causal bounds—i.e., rigorously identifying all possible answers that cannot be ruled out by an observed dataset. These results are made possible by the theoretical work proving that any causal problem can be transformed into an equivalent polynomial programming problem. Then, it becomes straightforward to derive a general algorithm for doing the problem reduction, which opens the door to a vast range of highly efficient polynomial programming solution techniques such as spatial branch-and-bound and linear programming relaxations. This method is an extension of the original Balke and Pearl (1997)’s linear programming technique and employs recent results in causality (Evans, 2018; Rosset et al., 2018).
@@
In the appendix, I have also derived bounds for estimands such as $\operatorname{\mathbb{E}}(Y_{a})$ and the ATE in two other scenarios: one with only factorial data, and another with both observational and factorial data, under the monotonicity assumption. There are notable limitations to those scenarios for cases involving treatments with more than two levels and other causal structures. However, these limitations can be addressed by employing the provided code, which makes use of autobounds.
Despite being bounded in scenarios with binary outcomes, these bounds can be extended to the discrete case efficiently. One simply needs to compute bounds for each possible outcome of $y$. For instance, if $Y$ has three levels $y_{0},y_{1},y_{2}$, one calculates the bounds three times: first for $\Pr(Y_{a}=y_{0})$, then for $\Pr(Y_{a}=y_{1})$, and $\Pr(Y_{a}=y_{2})$. During each bound computation, the other two levels are aggregated against the main one. When the outcome is continuous, a useful strategy suggested by both Balke and Pearl (1997) and Levis et al. (2023) is to calculate causal bounds for $\Pr(Y_{a}<k)$, for all $k$ within the bounded range of $Y$. In practice, this approach implies some form of discretization, ensuring valid bounds even if they are no longer sharp.
@@ 3.2 Sensitivity Analysis
The concept behind this sensitivity analysis acknowledges the strong nature of the AS assumption. Instead of outright accepting this, one could restrict the proportion of units exhibiting some level of interaction. For example, a parameter like $\theta=0.2$ acknowledges the potential presence of interactive units while also indicating that they may not exceed 20% of the population. Additionally, one can determine the maximum proportion of interactive units that could nullify the ATE or push it towards a specific value. Therefore, this procedure can be applied when presenting results on single-treatment effects, particularly with factorial data.